# Regression-Based Single-Point Zeroth-Order Optimization

### v1

@@ 3.1 Linear Regression-Based Single-Point Zeroth-Order Optimization (L-RESZO)
$\hat{X}_{t}\coloneq\begin{bmatrix}{\hat{\bm{x}}_{t}}^{\top}&1\\ {\hat{\bm{x}}_{t-1}}^{\top}&1\\ \vdots&\vdots\\ {\hat{\bm{x}}_{t-m+1}}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(d+1)},\quad\hat{\bm{y}}_{t}\coloneq\begin{bmatrix}f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-1})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})\end{bmatrix}\in\mathbb{R}^{m}.$ (11)
Consequently, when moving from iteration $t$ to $t+1$, the matrix $\hat{X}_{t+1}$ is actually formed by dropping the oldest sample point $\hat{\bm{x}}_{t-m+1}$ from $\hat{X}_{t}$ and adding the latest sample point $\hat{\bm{x}}_{t+1}$ to it. This leads to the update rule (12):
$\hat{X}_{t+1}^{\top}\hat{X}_{t+1}=\hat{X}_{t}^{\top}\hat{X}_{t}-\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}+\hat{\bm{x}}_{t+1}\hat{\bm{x}}_{t+1}^{\top}.$ (12)
Hence, given the available matrix inverse $(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ obtained in the last iteration, the new matrix inverse $(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}$ can be computed efficiently as (13) by using the Sherman–Morrison–Woodbury formula [21] twice in a row: one for dropping $\hat{\bm{x}}_{t-m+1}$ and the other for adding $\hat{\bm{x}}_{t+1}$;
$B=A+\frac{A\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}A}{1-\hat{\bm{x}}_{t-m+1}^{\top}A\hat{\bm{x}}_{t-m+1}},\qquad(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}=B-\frac{B\hat{\bm{x}}_{t+1}\hat{\bm{x}}_{t+1}^{\top}B}{1+\hat{\bm{x}}_{t+1}^{\top}B\hat{\bm{x}}_{t+1}},$ (13)
where $A\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ and $B\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t}-\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top})^{-1}$.
@@
(Role of Perturbation). From (11), it is seen that the rows of $\hat{X}_{t}$ are constructed by ${\hat{\bm{x}}_{k}}=\bm{x}_{k}\!+\!\delta\bm{u}_{k}$ for $k=t,\cdots,t\!-\!m\!+\!1$. Intuitively, adding the random perturbations $\delta\bm{u}_{k}$ promotes the full row rank property of $\hat{X}_{t}$, thus making $\hat{X}_{t+1}^{\top}\hat{X}_{t+1}$ tend to be invertible. This enhances the algorithmic stability of the RESZO method. The magnitude of perturbation is controlled by the value of the smoothing radius parameter $\delta$. Since the RESZO method essentially estimates the gradient at the perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, a smaller $\delta$ generally results in lower gradient estimation bias and promotes convergence precision. Accordingly, rather than selecting a fixed $\delta$, we propose an adaptive $\delta$ scheme for the RESZO algorithm in Section 4. The impacts and necessity of incorporating perturbations are further illustrated through numerical experiments in Section 5.2.
@@ 3.2 Quadratic Regression-Based Single-Point Zeroth-Order Optimization (Q-RESZO)
In addition to fitting a linear model, other more sophisticated surrogate models, such as a quadratic function, can be constructed based on available historical function evaluations to further enhance gradient estimation. At each iteration $t$, based on the second-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as:
$f(\hat{\bm{x}}_{k})\!=\!f(\hat{\bm{x}}_{t})\!+\!\nabla f(\hat{\bm{x}}_{t})^{\top}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})\!+\!\frac{1}{2}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})^{\top}H_{f}(\hat{\bm{x}}_{t})(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})\!+\!\epsilon_{t,k},\ \ k\!=\!0,1,\cdots,t,$ (14)
@@
where $H_{f}(\cdot)$ denotes the Hessian matrix of function $f$, and $\epsilon_{t,k}$ denotes the corresponding higher-order residual term. Accordingly, we can fit a quadratic surrogate function $f^{\mathrm{s2}}_{t}(\bm{x})$ (15) to locally approximate the original function $f(\bm{x})$ in (1):
$f^{\mathrm{s2}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}({\bm{x}}-\hat{\bm{x}}_{t})+\frac{1}{2}({\bm{x}}-\hat{\bm{x}}_{t})^{\top}\text{diag}(\bm{h}_{t})({\bm{x}}-\hat{\bm{x}}_{t})+c_{t}.$ (15)
@@
Here, we use the diagonal matrix $\text{diag}(\bm{h}_{t})$ with vector $\bm{h}_{t}\in\mathbb{R}^{d}$ to approximate the Hessian matrix, which reduces the number of variables for fitting. Similarly, the recursive linear regression in a sliding window is employed, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the quadratic surrogate model at each iteration $t$.
@@
Then, we solve the following least squares problem (16) to obtain the coefficients $(\bm{g}_{t},\bm{h}_{t},c_{t})$:
$\min_{\bm{g}_{t},\bm{h}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{[}f^{\mathrm{s2}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big{]}^{2}\Longrightarrow\begin{bmatrix}\bm{g}_{t}\\ \bm{h}_{t}\\ c_{t}\end{bmatrix}=(\tilde{X}_{t}^{\top}\tilde{X}_{t})^{\dagger}\tilde{X}_{t}^{\top}{\bm{y}}_{t},$ (16)
@@
where $\odot$ denotes the element-wise multiplication of two column vectors, and
@@
$\tilde{X}_{t}\coloneq\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,1}}\odot\bm{\Delta}^{x}_{t,1}\Big{)}^{\top}&1\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,2}}\odot\bm{\Delta}^{x}_{t,2}\Big{)}^{\top}&1\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,m}}\odot\bm{\Delta}^{x}_{t,m}\Big{)}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(2d+1)},\quad{\bm{y}}_{t}\coloneq\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m}\end{bmatrix}\in\mathbb{R}^{m}.$ (17)
@@
Accordingly, our proposed Quadratic Regression-based Single-point Zeroth-order Optimization (Q-RESZO) method for solving problem (1) follows the same framework as the L-RESZO method (i.e., Algorithm 1), except that it solves for $(\bm{g}_{t},\bm{h}_{t},c_{t})$ via (16) in Step 7 and performs the iterative update by (18) with the gradient estimator $\nabla f^{\mathrm{s2}}_{t}(\bm{x})$ in Step 8.
$\bm{x}_{t+1}=\bm{x}_{t}-\eta\,\nabla f^{\mathrm{s2}}_{t}(\bm{x}_{t})=\bm{x}_{t}-\eta\,\big{(}\bm{g}_{t}-\delta\bm{h}_{t}\odot\bm{u}_{t}\big{)}.$ (18)
@@
Additionally, the RESZO algorithms presented above are designed for solving unconstrained optimization problems as formulated in (1); nevertheless, they can be readily extended to constrained optimization problems using projection techniques [22] and primal-dual algorithms [23].
@@ 4 Theoretical Convergence Analysis
This section provides non-asymptotic convergence analysis, and we focus on the L-RESZO method. Throughout the analysis, we utilize the notation $[N]:=\{1,2,\dots,N\}$. We also denote $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$ to represent the gradient estimation error at an iteration $t$.
Throughout our analysis, we use the following standard assumption on the smoothness of the function $f$.
@@
For convenience, throughout the analysis, we will study the case where we do not include an intercept $c_{t}$. In this case, the approximate gradient $\bm{g}_{t}$ is the solution to the equation (19):
$\ \bm{g}_{t}=(S_{t}S_{t}^{\top})^{\dagger}S_{t}\Delta_{t}^{f},\quad S_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t})^{\top}\\ \end{bmatrix},\quad\Delta_{t}^{f}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-2})-f(\hat{\bm{x}}_{t})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-f(\hat{\bm{x}}_{t})\\ \end{bmatrix},$ (19)
where $S_{t}^{\top}\in\mathbb{R}^{(m-1)\times d},\Delta_{t}^{f}\in\mathbb{R}^{m-1}$.
Inspired by our theoretical analysis, we propose an adaptive $\delta$ scheme, given by (20), to enhance convergence precision:
[…]
where the smoothing radius parameter $\delta_{t}$ at each iteration $t$ is adaptively adjusted based on the norm of the last gradient estimate $\bm{g}_{t-1}$.
In our analysis, we will study the L-RESZO algorithm that employs this adaptive $\delta$ scheme, and numerical experiments on this scheme are performed in Section 5.2. We also make the following assumption on the regression procedure.
@@
Suppose $\bm{g}_{t}$ is as defined in (19). We assume that for any $t\geq m$,
$\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|\leq C_{d}\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$ (21)
for some constant $C_{d}>0$ depending on the dimension $d$ and smoothing radius sequence $\{\delta_{t}\}_{t}$.
Assumption 2 bounds the error of the approximate gradient by the distance between points used in the regression $\hat{\bm{x}}_{t-i}$ and the current point $\hat{\bm{x}}$; for clarity, we note that the error $\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})$ that appears in (21) is distinct from $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, since the former is the bias compared to the gradient at the perturbed point while the latter is the bias compared to the gradient at the actual iterate of the algorithm. While this assumption may appear unmotivated, we show in Appendix A.1 that it holds for $L$-smooth functions when $d=1$ with $C_{d}=1$. For general cases of $d$, Appendix C presents empirical evidence suggesting that the factor $C_{d}$ in Assumption 2 scales in the order of $\mathcal{O}(\sqrt{d})$ under the adaptive $\delta_{t}$ scheme in (20). While establishing a theoretical bound for $C_{d}$ is desired, it remains very challenging due to the technical difficulties discussed in Appendix A.1.1.
Before deriving our theoretical convergence results, we first specify the following condition on the initial $\tilde{\eta}$, the step-size during the initial update phase (before regression), which states that it should be sufficiently small in a certain sense. Our results later will all require $\tilde{\eta}$ to satisfy this condition.
@@
Recall that $\tilde{\eta}$ is the step-size of RESZO during the initial update phase (before regression). We pick $\tilde{\eta}$ such that for any $t\in\{0,1,\dots,m-1\}$, it satisfies [^1] We note that this is a relatively mild condition, and that we expect this to hold for any $\tilde{\eta}$ sufficiently small.
@@
We proceed with introducing our theoretical results. Our analysis is underpinned by the following key technical result, which bounds the gradient approximation error in terms of the norm of the true gradient. To be precise, we show that for any $10<c<$ 1, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
@@
Suppose that Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, for any $10<c<$ 1, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we can ensure that
$\|\xi_{t}\|=\|\bm{g}_{t}-\nabla f(\bm{x}_{t})\|\leq c\|\nabla f(\bm{x}_{t})\|$
@@
We provide a brief proof sketch here, deferring proof details to Appendix A.2. The proof consists of three key steps:
  1. 1.
    First, utilizing Assumption 2 and the adaptive smoothing radius selection, we can bound the gradient estimation error $\xi_{t}:=\bm{g}_{t}-\nabla f(x_{t})$ as
    $\|\xi_{t}\|\leq\mathcal{O}(C_{d}L)\sum_{j=1}^{m}\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|,$ (23)
    i.e. relate the gradient estimation error to a sum of the increments in the past $m$ updates.
  2. 2.
    Next, via an inductive argument (utilizing our condition on the initial step-size $\tilde{\eta}$), for any $t\geq m$ and $j\in[m]$, we show that we can always bound the one-step increment $\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$ as
    $\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|\leq 2\eta\|\nabla f(\bm{x}_{t-j})\|,$ (24)
    i.e. the size of each update can be bounded in terms of the step-size times the norm of the gradient at that step.
  3. 3.
    Utilizing a simple technical result (Lemma 5 in the Appendix A.2), we can show that whenever (24) holds, with an appropriate choice of $\eta$, the gradient norm changes only by an $O(1)$ multiplicative constant across $m$ iterations. Then, combining this fact with (23) and (24), we have that
    $\|\xi_{t}\|\leq$ $\ \mathcal{O}(C_{d}L)\sum_{j=1}^{m}\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
    $\leq$ $\ \mathcal{O}(C_{d}L)\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t-j})\|$
    $\leq$ $\ \mathcal{O}(C_{d}L)\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t})\|.$
    Using the condition that $m=\mathcal{O}(d)$, it follows then that for any $10<c<$ 1, by selecting $\eta=\frac{C}{dC_{d}L}$ for an appropriately small constant $C>0$, we can ensure that
    $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|,$
    as desired.
With Lemma 1 in hand, we can guarantee the following function value improvement per step.
@@
Suppose that Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we can ensure that
$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq-\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}$

### v2

@@ 3.1 Linear Regression-Based Single-Point Zeroth-Order Optimization (L-ReSZO)
$\hat{X}_{t}\!\coloneq\!\!\!\begin{bmatrix}{\hat{\bm{x}}_{t-1}}^{\top}\\ {\hat{\bm{x}}_{t-2}}^{\top}\\ \vdots\\ {\hat{\bm{x}}_{t-m+1}}^{\top}\end{bmatrix}\!\!\in\!\mathbb{R}^{(m-\!1)\!\times\!d},\ \hat{\bm{y}}_{t}\!\coloneq\!\!\!\begin{bmatrix}f(\hat{\bm{x}}_{t-1})\\ f(\hat{\bm{x}}_{t-2})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})\end{bmatrix}\!\!\in\!\mathbb{R}^{m-1}.$ (12)
As a result, when moving from iteration $t$ to $t+1$, the matrix $\hat{X}_{t+1}$ is actually formed by dropping the oldest sample point $\hat{\bm{x}}_{t-m+1}$ from $\hat{X}_{t}$ and adding the latest sample point $\hat{\bm{x}}_{t}$ to it. This leads to the update rule (13):
$\hat{X}_{t+1}^{\top}\!\hat{X}_{t+1}\!=\!\hat{X}_{t}^{\top}\!\hat{X}_{t}\!-\!\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}\!+\!\hat{\bm{x}}_{t}\hat{\bm{x}}_{t}^{\top}.$ (13)
Given the available matrix inverse $(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ obtained in the last iteration, the new matrix inverse $(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}$ can be computed efficiently as (14) by using the Sherman–Morrison–Woodbury formula [11] twice: one for dropping $\hat{\bm{x}}_{t-m+1}$ and the other for adding $\hat{\bm{x}}_{t}$;
$B=A+\frac{A\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}A}{1-\hat{\bm{x}}_{t-m+1}^{\top}A\hat{\bm{x}}_{t-m+1}},$ (14a)
$(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}=B-\frac{B\hat{\bm{x}}_{t}\hat{\bm{x}}_{t}^{\top}B}{1+\hat{\bm{x}}_{t}^{\top}B\hat{\bm{x}}_{t}},$ (14b)
$A\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1},\ B\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t}\!-\!\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top})^{-1}.$
@@
(Role of Perturbation). In (12), the rows of $\hat{X}_{t}$ are constructed by ${\hat{\bm{x}}_{k}}=\bm{x}_{k}\!+\!\delta_{k}\bm{u}_{k}$ for $k=t,\cdots,t\!-\!m\!+\!1$. Intuitively, adding random perturbations $\delta_{k}\bm{u}_{k}$ promotes the full row rank property of $\hat{X}_{t}$, thus making $\hat{X}_{t+1}^{\top}\hat{X}_{t+1}$ tend to be invertible. This is critical for ensuring the algorithmic stability of the ReSZO method. The magnitude of perturbations is controlled by the smoothing radius parameter $\delta_{k}$. In general, a smaller $\delta$ reduces the bias in gradient estimation and improves convergence precision, whereas a larger $\delta$ promotes greater exploration during gradient estimation. The proposed adaptive smoothing radius scheme (10) effectively balances these two aspects based on the current optimization state. The necessity of introducing perturbations and the effectiveness of this adaptive scheme are further demonstrated through numerical experiments in Section 5.2.
@@ 3.2 Quadratic Regression-Based Single-Point Zeroth-Order Optimization (Q-ReSZO)
In addition to fitting a linear model, other more sophisticated surrogate models, such as a quadratic function, can be constructed based on available historical function evaluations to further enhance gradient estimation. At each iteration $t$, based on the second-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as: for all $k\!=\!0,1,\cdots,t$,
$f(\hat{\bm{x}}_{k})=f(\hat{\bm{x}}_{t})\!+\!\nabla f(\hat{\bm{x}}_{t})^{\top}\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)+\frac{1}{2}\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)^{\top}H_{f}(\hat{\bm{x}}_{t})\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)\!+\!\epsilon_{t,k}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}),$ (15)
@@
where $H_{f}(\cdot)$ is the Hessian matrix of function $f$, and $\epsilon_{t,k}(\cdot)$ denotes the corresponding higher-order residual term. Accordingly, we can fit a quadratic surrogate function $f^{\mathrm{s2}}_{t}(\bm{x})$ (16) to locally approximate the original function $f$ in (1):
$\ f^{\mathrm{s2}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}\big({\bm{x}}-\hat{\bm{x}}_{t}\big)+\frac{1}{2}\big({\bm{x}}-\hat{\bm{x}}_{t}\big)^{\top}\text{diag}(\bm{h}_{t})\big({\bm{x}}-\hat{\bm{x}}_{t}\big).$ (16)
@@
Here, we use the diagonal matrix $\text{diag}(\bm{h}_{t})$ with vector $\bm{h}_{t}\in\mathbb{R}^{d}$ to approximate the Hessian matrix, which reduces the number of variables for fitting. Similarly, a constant intercept term is omitted from the quadratic surrogate function (16) to enforce $f^{\mathrm{s2}}_{t}(\hat{\bm{x}}_{t})=f(\hat{\bm{x}}_{t})$ and to mitigate model bias. In addition, the recursive linear regression in a sliding window is employed, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the quadratic surrogate model (16) at each iteration $t$.
@@
Then, we solve the following least squares problem (17) to obtain the coefficients $\{\bm{g}_{t},\bm{h}_{t}\}$:
$\min_{\bm{g}_{t},\bm{h}_{t}}\,\sum_{i=1}^{m}\Big[f^{\mathrm{s2}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big]^{2}\ \Longrightarrow\ \begin{bmatrix}\bm{g}_{t}\\ \bm{h}_{t}\end{bmatrix}=(\tilde{X}_{t}^{\top}\tilde{X}_{t})^{\dagger}\tilde{X}_{t}^{\top}{\bm{y}}_{t},$ (17)
@@
where $\odot$ denotes the element-wise multiplication of two column vectors, $\tilde{X}_{t}\in\mathbb{R}^{(m-1)\times 2d}$ and ${\bm{y}}_{t}\in\mathbb{R}^{m-1}$ with
@@
$\tilde{X}_{t}\!\coloneq\!\!\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&\frac{1}{2}\big({\bm{\Delta}^{x}_{t,1}}\!\odot\!\bm{\Delta}^{x}_{t,1}\big)^{\top}\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&\frac{1}{2}\big({\bm{\Delta}^{x}_{t,2}}\!\odot\!\bm{\Delta}^{x}_{t,2}\big)^{\top}\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m\!-\!1}}^{\top}&\frac{1}{2}\big({\bm{\Delta}^{x}_{t,m\!-\!1}}\!\odot\!\bm{\Delta}^{x}_{t,m\!-\!1}\big)^{\top}\end{bmatrix}\!,{\bm{y}}_{t}\!\coloneq\!\!\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m\!-\!1}\end{bmatrix}\!$. (18)
@@
Accordingly, our proposed Quadratic Regression-based Single-point Zeroth-order Optimization (Q-ReSZO) method for solving problem (1) follows the same framework as the L-ReSZO method (i.e., Algorithm 1), except that it solves for $\{\bm{g}_{t},\bm{h}_{t}\}$ via (17) in Step 8 and performs the iterative update by (19) with the gradient estimator $\nabla f^{\mathrm{s2}}_{t}(\bm{x})$ in Step 9.
$\bm{x}_{t+1}\!=\!\bm{x}_{t}\!-\!\eta\nabla f^{\mathrm{s2}}_{t}(\bm{x}_{t})\!=\!\bm{x}_{t}\!-\!\eta\big(\bm{g}_{t}-\delta_{t}\bm{h}_{t}\odot\bm{u}_{t}\big).$ (19)
@@
The ReSZO algorithms presented above aim for solving unconstrained optimization problems as formulated in (1); nevertheless, they can be readily extended to constrained optimization problems using projection techniques [6] and primal-dual algorithms [5].
@@ 4 Theoretical Convergence Analysis
This section presents the non-asymptotic convergence analysis for the L-ReSZO method. Throughout the analysis, we use the notation $N]\!\coloneqq\!\{1,2,\dots,N\}$, and $\xi_{t}\!\coloneqq\!\bm{g}_{t}\!-\!\nabla f(\bm{x}_{t})$ denotes the gradient estimation error at iteration $t$.
Throughout our analysis, we make the following standard assumption on the smoothness of the function $f$.
@@
We recall the approximate gradient $\bm{g}_{t}$ can be written as equation (20):
$\qquad\qquad\qquad\qquad\bm{g}_{t}=(S_{t}S_{t}^{\top})^{\dagger}S_{t}\Delta_{t}^{f},$ (20)
$S_{t}^{\top}\!:=\!\!\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}\!-\!\hat{\bm{x}}_{t})^{\top}\\ \end{bmatrix}\!,\Delta_{t}^{f}\!:=\!\!\begin{bmatrix}f(\hat{\bm{x}}_{t-1})\!-\!f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-2})\!-\!f(\hat{\bm{x}}_{t})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})\!-\!f(\hat{\bm{x}}_{t})\\ \end{bmatrix}$,
where $S_{t}^{\top}=X_{t}\in\mathbb{R}^{(m-1)\times d},\Delta_{t}^{f}=\bm{y}_{t}\in\mathbb{R}^{m-1}$.
We also make the following assumption on the regression procedure.
@@
Suppose that $\bm{g}_{t}$ is as defined in (20), and that $m>d$. We assume that for any $t\geq m$,
$\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|\leq C_{d,m}\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$ (21)
for some constant $C_{d,m}>0$ depending on the dimension $d$ and $m$. In particular, when the regression sample size is picked as $m=\mathcal{O}(d)$, we denote $C_{d}:=C_{d,m}$.
Assumption 2 bounds the error of the approximate gradient by the distance between points used in the regression $\hat{\bm{x}}_{t-i}$ and the current point $\hat{\bm{x}}_{t}$. For clarity, we note that the error $\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})$ that appears in (21) is distinct from $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, since the former is the bias compared to the gradient at the perturbed point while the latter is the bias compared to the gradient at the actual iterate of the algorithm. We note that this assumption is difficult to rigorously prove, since it relies on the intricate behavior of the pseudoinverse of the random matrix $S_{t}S_{t}^{\top}$. We provide heuristic reasoning for its validity in Appendix B.1, showing that $C_{d,m}$ scales at worst as $\mathcal{O}(\sqrt{d}{\sqrt{m}})=\mathcal{O}(d)$ (when we set $m=\mathcal{O}(d)$). In practice, we set $m=\mathcal{O}(d)$ in our algorithm and present empirical evidence in Appendix C suggesting that in this setting, the factor $C_{d}:=C_{d,m}$ in Assumption 2 actually scales on the order of $\mathcal{O}(\sqrt{d})$.
Before deriving our theoretical convergence results, we first specify the following condition on the initial $\tilde{\eta}$, namely the stepsize during the initial phase (before regression), indicating that it should be sufficiently small in a certain sense. Our subsequent results require that $\tilde{\eta}$ satisfies this condition.
@@
Recall that $\tilde{\eta}$ is the step-size of ReSZO during the initial update phase (before regression). We pick $\tilde{\eta}$ such that for any $t\in\{0,1,\dots,m-1\}$, it satisfies [^1] We note that this is a relatively mild condition and expect it to hold for any sufficiently small $\tilde{\eta}$.
@@
Our analysis is underpinned by the following key technical result, which bounds the gradient approximation error in terms of the norm of the true gradient. To be precise, we show that for any $0<c<1$, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
@@
Suppose that Assumptions 1 and 2 hold, and that $m>d$. Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied. Then, for any $0<c<1$, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, we have that
$\|\xi_{t}\|=\|\bm{g}_{t}-\nabla f(\bm{x}_{t})\|\leq c\|\nabla f(\bm{x}_{t})\|,$ (23)
@@
We defer the proof to Appendix B.
With Lemma 1 in hand, we can guarantee the following function value improvement per step; the proof of this result is provided in Appendix B.
@@
Suppose that Assumptions 1 and 2 hold, and that $m>d$. Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied. Then, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, we have that
$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq-\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2},$ (24)